A liquid in a lab has a density of 1.17 g/cm^3. What is the volume in liters of 3.02 kg of the liquid?


1
Expert's answer
2019-09-13T05:10:48-0400

"V = \frac{m}{\rho} = \frac{3.02\text{ kg}}{1.17\text{ g/cm}^3} = \frac{3.02 \cdot 10^3 \text{ g}}{1.17\text{ g}} \text{cm}^3 \approx 2.58 \cdot 10^3 \text{cm}^3 = 2.58\text{ L (liters)}."